4. Find the surface area of part of the paraboloid z =x2 + y2 cut of by the plane z = 4.
Answer #1

Given, z= x + y 2 cut by z = 4 when z=4 =) x+y+=4 so, 02122 And OLOL 271 az 02 2x дz by pour = 2y. Surface Area is given as S

ALL THE BEST ...,!